Find an example of Python program on the Web and provide a link for the source.


Discuss what the program does and whether it has inputs, processes, and outputs.
Answer #1

Answer:

According to the policies, can't provide external link.

Here I am writing a basic python program with screenshots and output:


a = int(input('Enter first number: '))
b = int(input('Enter second number: '))

addition = a+b
difference = a-b
multiplication = a*b
division = a/b

print("Addition:", addition)
print("Difference:", difference)
print("Multiplication:", multiplication)
print("Division:", division)

Input: The program takes input from the user, 2 integers a and b

Processing: It computes the addition, subtraction, multiplication and division of those 2 variables and store them in the new variables.

Output: It prints all the variables created during processing.

Screenshot:

Output:

PLEASE UPVOTE IF YOU FOUND THIS HELPFUL!